Let r(t)=0. Then ∫r(t)dt is equal to
Select one:
a. ⟨0,t,t⟩+c where c is an arbitrary constant vector

b. ⟨t,t,0⟩+c where c is an arbitrary constant vector

c. c where c is an arbitrary constant vector

d. ⟨t,t,t⟩+c where c is an arbitrary constant vector

e. ⟨t,0,t⟩+c where c is an arbitrary constant vector

f. ⟨t,0,0⟩+c where c is an arbitrary constant vector
1
Expert's answer
2020-05-19T19:12:01-0400

the answer is (c)

Since position vector "r(t)=x(t) i+y(t)j+z(t)k"

and from the hypothesis that "r(t)=0=0 i+0j+0k"

Therefore "\int r(t)dt=\int( 0 i+0j+0k )dt"

and then

"\int r(t)dt=c_{1} i+c_{2}j+c_{3}k=c,"

where "c" is an arbitrary constant vector